OpenStudy (hihi67):

For the form of the equation ax^2+bx+c=0 you know c (-7.2) and both answers x1 = -2 x2 = 4 what are then a and b?

OpenStudy (hihi67):

@ash2326 @Callisto @dmac83 @experimentX

OpenStudy (hihi67):

@jhonyy9

OpenStudy (hihi67):

@sarahusher @kirad @LostInSchool

OpenStudy (hihi67):

@whambamsham @worldcoin @eliassaab @robtobey @RANE

OpenStudy (hihi67):

@Triton

OpenStudy (hihi67):

@theycallmerhea

OpenStudy (hihi67):

@Orion1213

OpenStudy (hihi67):

@phi

OpenStudy (hihi67):

Does anybody think they can help?

OpenStudy (phi):

Try this (as step 1): in your equation \[ ax^2+bx+c=0 \] replace c with -7.2 and x with -2 and simplify as much as possible what do you get ?

OpenStudy (hihi67):

Nooo that's not what I meant sorry.

OpenStudy (hihi67):

x1 and x2 are the answers...

OpenStudy (hihi67):

Oh, lol yeah I get what you said now lol

OpenStudy (hihi67):

But that's not how we're supposed to solve it.

OpenStudy (phi):

usually we solve for "x". (and we know a,b, and c) but in this case, we are going to solve for a and b, (but we know x) so replace c with -7.2 and x with -2 and simplify as much as possible what do you get ?

OpenStudy (hihi67):

Okay just a minute.

OpenStudy (hihi67):

4a + -2b - 7.2 = 0

OpenStudy (phi):

looks good. now do the same thing, but use x= 4 we will get another equation

OpenStudy (hihi67):

Okay... 16a + 4b - 7.2 = 0

OpenStudy (phi):

yes, now we have 2 equations and 2 unknowns. There are ways to solve these things. 16a + 4b - 7.2 = 0 4a + -2b - 7.2 = 0 one way is multiply the bottom equation by 2 (that means both sides all terms) then add this "new" equation to the first equation.

OpenStudy (hihi67):

Maybe, but I'm sorry cause we didn't learn ANY of this.. we did the thing with Vieta though, and I'm pretty sure that's how we should solve this.

OpenStudy (hihi67):

Do you know what I'm talking about?

OpenStudy (phi):

Yes, you can do it that way. x1 + x2 = -b/a x1 * x2= c/a I would use the 2nd equation first, to solve for a. can you do that ?

OpenStudy (hihi67):

Okay, that's what I thought, thanks!

OpenStudy (hihi67):

Wait, so what is -b/a exactly?

OpenStudy (phi):

a , b and c are the coefficients in front of a x^2 + bx + c = 0 you need to find a and b in the 2nd equation x1 * x2 = c/a they give you x1 and x2 and c that leaves a the only unknown. replace the letters with the numbers they gave you, and simplify what do you get ?

OpenStudy (hihi67):

Yeah I know, but how will knowing -b/a help me??

OpenStudy (phi):

It will be clear in a moment. First, do x1 * x2 = c/a

OpenStudy (hihi67):

Okay, -2 times 4 is -8

OpenStudy (hihi67):

And since c is 7.2

OpenStudy (phi):

be careful, c= -7.2

OpenStudy (hihi67):

a is .9

OpenStudy (hihi67):

Okay, it sucks cause I did exactly this, I just made a calculating error before lol But thanks a lot! :D I can do the rest by myself!!

OpenStudy (phi):

yes, a is 0.9 now that you know a, the first equation x1+x2 = -b/a has only *1 unknown*, the "b", so we can solve for it